OpenStudy (anonymous):

so so very lost with this. Find the reference angle for the given angle. (a) 10π/3 (b) 33π/4 (c) −35π/6 I would like to understand how to do one of them at least.

OpenStudy (anonymous):

just subtract multiples of 2pi so that 0<the resultant < 2pi

OpenStudy (anonymous):

or add, if original is negative

OpenStudy (anonymous):

so for A) i got \[-4\pi/3\]

OpenStudy (anonymous):

but that was the wrong answer

OpenStudy (anonymous):

i did that for A, B, and C. but i got the wrong answers

OpenStudy (anonymous):

is \[ 0<\frac {-4 \pi}{3}< 2\pi\]

OpenStudy (psymon):